OpenStudy (anonymous):

Last Wednesday it snowed 3 1/4 in. On Sunday it snowed 3 4/5 in. On which day did it snow more? How much more?

OpenStudy (jhannybean):

Convert 3 1/4 to a whole fraction: \(\frac{4 \cdot 3 +1}{4} = \frac{13}{4} \ in.\) Convert 3 4/5 into a whole fraction: \(\frac{5 \cdot 3 +4}{5} = \frac{19}{5}\ in.\) Write these both : \(\frac{13}{4} = 3.25\ in.\) and \(\frac{19}{5} = 3.8 \ in.\)

OpenStudy (jhannybean):

So Wednesday = 3.25 in , and Sunday would = 3.8 in. Analyzing these two, which one seems bigger? To find "how much more", just subtract the two to find the difference between them. 3.8 - 3.25 = ?